Home - Search - Browse - Alphabetic Index: 0- 1- 2- 3- 4- 5- 6- 7- 8- 9
A- B- C- D- E- F- G- H- I- J- K- L- M- N- O- P- Q- R- S- T- U- V- W- X- Y- Z
More Details for 2008-06-20
ISS On-Orbit Status 06/20/08

Crew wake-up was half an hour earlier (1:30am EDT) for the Orlan training exercise, enabling suit telemetry over RGS (Russian Groundsites).

Sleep time tonight will be back at regular time (5:30pm).

In the DC1 'Pirs', the CDR & FE-1 continued preparations for the EVA-20 on 7/10-11, with the DC1 comm links configured for their presence. Activities by Volkov & Kononenko today focused on

Functional testing of the Russian BETA-08 ECG (electrocardiogram) lead cable belts, worn under the Orlan-M suits, using the Gamma-1M medical complex from the PKO medical exam panel,
Transmission tests, after setting up the Orlan 'Tranzit' communications links via the suits' BRTA radio telemetry units, with the ground via RGS VHF, for spacesuit voice, telemetry & biomedical parameters; followed by
Restoring nominal communications setup in the DC1, and
Configuring the backup Orlan #25 for stowage mode, leaving Orlans #26 (FE-1) & #27 (CDR) ready for use on the upcoming spacewalk.
In the U.S. 'Quest' Airlock, FE-2 Chamitoff continued post-1J EVA activities, terminating the 13-hr discharge process on the last 16-volt EMU (Extravehicular Mobility Unit) battery in the BCM-3 (Battery Charger Module 3) from yesterday. (BCM-4 is currently nonfunctional. The discharge process, originally handled manually by a crewmember, is an automated procedure controlled from an A31p SSC (Station Support Computer) laptop with a special DOS application.)

The CDR serviced the Russian BMP (Harmful Impurities Removal System), starting the "bake-out"-to-vacuum cycle on absorbent bed #2 of the regenerable dual-channel filtration system. The regen process will be terminated tonight at ~5:15pm EDT. Filter bed #1 was regenerated yesterday. (Regeneration of each of the two cartridges takes about 12 hours and is conducted only during crew awake periods. The BMP's regeneration cycle is regularly done every 20 days.)

After breaking out the auditory test equipment, Gregory Chamitoff took the periodic O-OHA (on-orbit hearing assessment) test, a 30-min. NASA environmental health systems examination to assess the efficacy of acoustic countermeasures, using a special MEC (Medical Equipment Computer) laptop application. It was his first O-OHA session. (The O-OHA audiography test involves minimum audibility measurements for each ear over a wide range of frequencies (0.25-10 kHz) and sound pressure levels, with the crewmembers using individual-specific Prophonics earphones, Bose ANC headsets and the SLM (sound level meter). To conduct the testing, the experimenter is supported by special EarQ software on the MEC, featuring an up/down-arrow-operated slider for each test frequency that the crewmember moves to the lowest sound pressure level at which the tone can still be heard. The baseline test is required not later than about Flight Day 14 for each new Expedition and is then generally performed once per month. Note: There have been temporary hearing deficits documented on some U.S. and Russian crewmembers, all of which recovered to pre-mission levels.)

Later, after setting up the video equipment, Chamitoff underwent his first PFE (Periodic Fitness Evaluation) session on the CEVIS cycle ergometer, with Volkov assisting in filming Greg and obtaining his blood pressure measurements during the CEVIS workout. Subsequently, the FE-2 completed data entry and stowed the PFE and video hardware. (The U.S. PFE w/o Blood Lab is a monthly 1.5-hr. procedure which checks up on blood pressure and electrocardiogram (ECG) during programmed exercise on the CEVIS (Cycle Ergometer with Vibration Isolation) in the Lab. Readings are taken with the BP/ECG (blood pressure/electrocardiograph) and the HRM (heart rate monitor) watch with its radio transmitter. BP/ECG provides automated noninvasive systolic and diastolic blood pressure measurements while also monitoring and displaying accurate heart rates on a continual basis at rest and during exercise.)

Continuing Kibo laboratory outfitting, the FE-2 worked in the JPM (JEM Pressurized Module), removing the launch lock on the JPM internal port TV camera.

For setting up the Kibo SLT (Station Laptop Terminal), Chamitoff powered off the SSC-14 (Station Support Computer 14), relocated it in the JPM and plugged it into a new UOP (Utility Outlet Panel).

In the SM (Service Module), Kononenko downloaded protocol/log files of the BRI Smart Switch Router to the RSS1 laptop for subsequent dumping to the ground via OCA, for specialists to review issues, then rebooted the RRS1. (BRI is part of the RS OpsLAN network with connections to the three SSC clients, the Ethernet tie-in with the US network, and a network printer in the RS (Russian segment).)

The FE-1 also performed functional verification of the ongoing radiation data monitoring for flux & dose rate data with the Matryoshka-R radiation payload via its Lulin-5 electronics box, supported by specialist tagup via S-band. (Accumulated readings were recorded on a log sheet for subsequent downlink to TsUP/Moscow over the BSR-TM payload data channel.)

Later, Kononenko unstowed the Lulin-ISS radiation complex kit, set up its ICU (Interface Control Unit, Russian: BUI) and connected it to a power outlet (PPS-26) with four PILLE radiation dosimeters, to allow their battery to recharge for the next 24 hrs. (Done before on 4/24.)

Performing routine maintenance on the primary CSA-CP (Compound Specific Analyzer-Combustion Products) instrument, Gregory replaced its battery (#1080) with a fresh one (#1329), then re-deployed the unit.

Kononenko conducted the periodic (currently daily) checkout/verification of IP-1 airflow sensors in the various RS hatchways, including the passageways PrK (SM Transfer Compartment) -ATV, PrK -RO (SM Working Compartment), PkhO (SM Transfer Tunnel) -RO, PkhO -DC1, PkhO -FGB PGO, FGB PGO -FGB GA, FGB GA -Node-1.

Oleg also set up the pumping equipment and initiated (later closed out) the periodic transfer of urine from 9 EDV-U containers in the SM to the Rodnik BV2 tank of Progress M-64/29P, then flushed the system with ~5 L of disinfectant solution. (Transfer hoses were discarded as usual.)

As part of regular preventive maintenance of RS ventilation systems, Sergey meanwhile applied vacuum cleaner and soft brush to cleaning the detachable VT7 fan screens 1, 2 & 3 of the three SOTR (Thermal Control System) gas-liquid heat exchangers (GZhT4) In the FGB (Funktsionalnyi-Grusovoi Blok).

The FE-2 unstowed, activated and checked out a new AED (Automated External Defibrillator) in the CHeCS (Crew Health Care Systems) rack for future use on the ISS, disposing of the old defibrillator. (The AED is a portable electronic device that automatically diagnoses the potentially life threatening cardiac arrhythmias of ventricular fibrillation and ventricular tachycardia in a patient. It then can treat them through defibrillation, i.e., the application of electrical therapy which stops the arrhythmia, allowing the heart to reestablish an effective rhythm. AEDs are generally either held by trained personnel who will attend events or are public access units which can be found in places including corporate and government offices, shopping centers, airports, restaurants, casinos, hotels, sports stadiums, schools and universities, community centers, fitness centers, health clubs and any other location where people may congregate.)

Afterwards, Gregory deployed a new CCPK (Crew Contamination Protection Kit, #1002) and new Eyewash Kit in the SM (near the SVO-ZV water tap/port).

Chamitoff also deployed new medical SODF (Station Operations Data File) books and updated checklists (cue cards), easily accessible in a Lab rack drawer used for medical items (LAB1D4_D1).

The FE-2 took measurements for the regular atmospheric status check for ppCO2 (Carbon Dioxide partial pressure) in the Lab, SM (at panel 449) and COL (Columbus Orbital Laboratory), using the hand-held CDMK (CO2 Monitoring Kit) #1002). (The battery pack was to be replaced with the one from unit #1009 if necessary.)

With the U.S. CDRA (Carbon Dioxide Removal Assembly) deactivated by the ground this morning (~5:00am-9:45am) and its cooling no longer required, Chamitoff demated and took down the ITCS LTL (Internal Thermal Control System/Low Temperature Loop) jumper at the CDRA-supporting LAB1D6 rack.

After the hold put by the ground on the use of O2 bottles (due to some corrosion observed on O2 pressure gauges on ground units), Gregory took documentary close-up photography of the OUM-PFE (Oxygen Uptake Measurement - Periodic Fitness Evaluation) payload's DS (Gas Delivery System) for ground analysis.

The crew conducted their regular 2.5-hr. physical workout program (about half of which is used for setup & post-exercise personal hygiene) on the CEVIS cycle ergometer (FE-2/PFE), TVIS treadmill (CDR, FE-1), RED resistive exercise device (C DR, FE-1, FE-2) and VELO bike with bungee cord load trainer (CDR, FE-1).

Afterwards, Sergey transferred the exercise data file to the MEC (Medical Equipment Computer) laptop for downlink, including the daily wristband HRM (Heart Rate Monitor) data of the workouts on RED, followed by their erasure on the HRM storage medium (done six times a week).

Chamitoff conducted the periodic (every two weeks) inspection of the RED canister bolts, to be re-tightened if required. (Deferred from 6/17.)

The CDR completed the routine maintenance of the SOZh system (Environment Control & Life Support System, ECLSS) in the SM, including ASU toilet facilities systems/replaceables.

Volkov also performed the daily IMS (Inventory Management System) maintenance, updating/editing its standard 'delta file' including stowage locations, for the regular weekly automated export/import to its three databases on the ground (Houston, Moscow, Baikonur).

Still remaining on the voluntary task list item for Kononenko & Volkov today was an audit of expired Expedition 16 food rations, with repacking & preparation of food packages for disposal on the ATV. (To clear storage space for cargo items delivered on Progress M-64/29P.)

At ~5:10am EDT, the crew held the regular (nominally weekly) tagup with the Russian Flight Control Team (GOGU/Glavnaya operativnaya gruppa upravleniya = 'Main Operative Control Group'), including Shift Flight Director (SRP), at TsUP-Moscow via S-band/audio, phone-patched from Houston and Moscow.

At ~6:05am, Sergey & Oleg linked up with TsUP stowage specialists via S-band to conduct the weekly IMS tagup, discussing stowage issues, equipment locations and Progress cargo transfers. (Issues discussed today concerned Progress 29P unloading status, number of remaining ASU toilet replaceable rings, whereabouts of a specific cable, etc.)

At ~3:10pm, the ISS crew is scheduled for their regular weekly tagup with the Lead Flight Director at JSC/MCC-H via S-band/audio. (S/G-2 (Space-to-Ground 2) phone patch via SSC (Station Support Computer)).

MT Translation: The planned move of the MT (Mobile Transporter) from WS-6 (Worksite 6) to WS-4 is now scheduled for today at ~3:00pm-4:30pm. Russian RCS (Reaction Control System) thrusters will be disabled for this purpose at 3:00pm - 5:30pm.

WRM Update: An updated Water Recovery Management 'cue card' was uplinked for the crew's reference. (The new card (17-0002M) lists 36 CWCs (~1448.4 L total) for the four types of water identified on board: technical water (694.6 L, for Elektron, flushing, hygiene, incl. 553.4 L non-usable water because of Wautersia bacteria), potable water (706.7 L, incl. 260.6 L currently on hold), condensate water (39.2 L), waste/EMU dump and other (7.9 L). Wautersia bacteria are typical water-borne microorganisms that have been seen previously in ISS water sources. These isolates pose no threat to human health.)

CEO (Crew Earth Observations) photo targets uplinked for today were Mount Vesuvius (Vesuvius is of course famous for its 79 AD eruption destroying the towns of Pompeii and Herculaneum. The volcano is still considered active and could erupt in the near future. The volcano's proximity to Naples is of great concern. Trying for nadir views of the volcano), Mt. Etna (Mount Etna, towering above Catania, Sicily's second largest city, has one of the world's longest documented records of historical volcanism, dating back to 1500 BC. Historical lava flows of basaltic composition cover much of the surface of this massive volcano, whose edifice is the highest and most voluminous in Italy. Context views of the volcano were requested), Central-Arizona Phoenix (the Central-Arizona Phoenix site is part of the Long Term Ecological Research (LTER) suite of sites. The main objective for these sites is to document the land cover/land use change on a seasonal basis. Looking for land cover or land use boundaries and document with overlapping frames), and Teide Volcano (the large triangular island of Tenerife is composed of a complex of overlapping volcanoes that have remained active into historical time. The United Nations Committee for Disaster Mitigation have designated Teide as a Decade Volcano. It is considered to be the 13th most dangerous volcano in the world due to its proximity to several major towns and the nearby city of Puerto de la Cruz. Contest views of the volcano were requested).

Home - Search - Browse - Alphabetic Index: 0- 1- 2- 3- 4- 5- 6- 7- 8- 9
A- B- C- D- E- F- G- H- I- J- K- L- M- N- O- P- Q- R- S- T- U- V- W- X- Y- Z
© 1997-2017 Mark Wade - Contact
© / Conditions for Use